\begin{thebibliography}{}
%\cite{Maldacena}	
\bibitem{Maldacena}
J. M. Maldacena, 
%``The Large N limit of superconformal field theories and supergravity,''
 Int.\ J.\ Theor.\ Phys.\  {\bf 38} (1999) 1113 [Adv.\ Theor.\ Math.\ Phys.\  {\bf 2} (1998) 231] [hep-th/9711200].	
	
	
%\cite{Gubser:1998bc}
\bibitem{Gubser:1998bc}
S.~S.~Gubser, I.~R.~Klebanov and A.~M.~Polyakov,
%``Gauge theory correlators from noncritical string theory,''
Phys. Lett. B \textbf{428}, 105-114 (1998)
doi:10.1016/S0370-2693(98)00377-3
[arXiv:hep-th/9802109 [hep-th]].	


%\cite{Aharony:1999ti}
\bibitem{Aharony:1999ti}
O.~Aharony, S.~S.~Gubser, J.~M.~Maldacena, H.~Ooguri and Y.~Oz,
%``Large N field theories, string theory and gravity,''
Phys. Rept. \textbf{323} (2000), 183-386
doi:10.1016/S0370-1573(99)00083-6
[arXiv:hep-th/9905111 [hep-th]].


	
%\cite{Witten:1998qj}
\bibitem{Witten:1998qj}
E.~Witten,
%``Anti-de Sitter space and holography,''
Adv. Theor. Math. Phys. \textbf{2}, 253-291 (1998)
doi:10.4310/ATMP.1998.v2.n2.a2
[arXiv:hep-th/9802150 [hep-th]].

%\cite{Jiang:2017imk}
\bibitem{Jiang:2017imk}
W.~J.~Jiang, H.~S.~Liu, H.~Lu and C.~N.~Pope,
%``DC Conductivities with Momentum Dissipation in Horndeski Theories,''
JHEP \textbf{07} (2017), 084
doi:10.1007/JHEP07(2017)084
[arXiv:1703.00922 [hep-th]].

%\cite{Baggioli:2017ojd}
\bibitem{Baggioli:2017ojd}
M.~Baggioli and W.~J.~Li,
%``Diffusivities bounds and chaos in holographic Horndeski theories,''
JHEP \textbf{07} (2017), 055
doi:10.1007/JHEP07(2017)055
[arXiv:1705.01766 [hep-th]].



%\cite{Shuryak:2003xe}
\bibitem{Shuryak:2003xe}
E.~Shuryak,
%``Why does the quark gluon plasma at RHIC behave as a nearly ideal fluid?,''
Prog. Part. Nucl. Phys. \textbf{53}, 273-303 (2004)
doi:10.1016/j.ppnp.2004.02.025
[arXiv:hep-ph/0312227 [hep-ph]].
 
 
 %\cite{Hartnoll:2009sz}
 \bibitem{Hartnoll:2009sz}
 S.~A.~Hartnoll,
 %``Lectures on holographic methods for condensed matter physics,''
 Class. Quant. Grav. \textbf{26} (2009), 224002
 doi:10.1088/0264-9381/26/22/224002
 [arXiv:0903.3246 [hep-th]].
 
%\cite{Peet:1998wn}
%\bibitem{Peet:1998wn}
%A.~W.~Peet and J.~Polchinski,
%``UV / IR relations in AdS dynamics,''
%Phys. Rev. D \textbf{59}, 065011 (1999)
%doi:10.1103/PhysRevD.59.065011
%[arXiv:hep-th/9809022 [hep-th]]. 


\bibitem{D.T. Son}
D.T. Son, % Hydrodynamics and gauge/gravity duality,
 Nuclear Physics B (Proc. Suppl.) 192--193 (2009) 113--118.



\bibitem{Policastro}
G.~Policastro, D.~T.~Son and A.~O.~Starinets,%``From AdS/CFT correspondence to hydrodynamics,''
JHEP {\bf 0209}, 043 (2002)
[hep-th/0205052].




%\cite{Son:2007vk}
\bibitem{Son:2007vk}
D.~T.~Son and A.~O.~Starinets,
%``Viscosity, Black Holes, and Quantum Field Theory,''
Ann. Rev. Nucl. Part. Sci. \textbf{57} (2007), 95-118
doi:10.1146/annurev.nucl.57.090506.123120
[arXiv:0704.0240 [hep-th]].



%\cite{Kovtun:2004de}
\bibitem{Kovtun:2004de}
P.~Kovtun, D.~T.~Son and A.~O.~Starinets,
%``Viscosity in strongly interacting quantum field theories from black hole physics,''
Phys. Rev. Lett. \textbf{94} (2005), 111601
doi:10.1103/PhysRevLett.94.111601
[arXiv:hep-th/0405231 [hep-th]].


%\cite{Bhattacharyya:2007vjd}
\bibitem{Bhattacharyya:2007vjd}
S.~Bhattacharyya, V.~E.~Hubeny, S.~Minwalla and M.~Rangamani,
%``Nonlinear Fluid Dynamics from Gravity,''
JHEP \textbf{02} (2008), 045
doi:10.1088/1126-6708/2008/02/045
[arXiv:0712.2456 [hep-th]].


%\cite{Policastro:2001yc}
\bibitem{Policastro:2001yc}
G.~Policastro, D.~T.~Son and A.~O.~Starinets,
%``The Shear viscosity of strongly coupled N=4 supersymmetric Yang-Mills plasma,''
Phys. Rev. Lett. \textbf{87} (2001), 081601
doi:10.1103/PhysRevLett.87.081601
[arXiv:hep-th/0104066 [hep-th]].


%\cite{Rangamani:2009xk}
\bibitem{Rangamani:2009xk}
M.~Rangamani,
%``Gravity and Hydrodynamics: Lectures on the fluid-gravity correspondence,''
Class. Quant. Grav. \textbf{26} (2009), 224003
doi:10.1088/0264-9381/26/22/224003
[arXiv:0905.4352 [hep-th]].



%\cite{Bhattacharya:2011tra}
\bibitem{Bhattacharya:2011tra}
J.~Bhattacharya, S.~Bhattacharyya, S.~Minwalla and A.~Yarom,
%``A Theory of first order dissipative superfluid dynamics,''
JHEP \textbf{05} (2014), 147
doi:10.1007/JHEP05(2014)147
[arXiv:1105.3733 [hep-th]].







%\cite{Kovtun:2012rj}
\bibitem{Kovtun:2012rj}
P.~Kovtun,
%``Lectures on hydrodynamic fluctuations in relativistic theories,''
J. Phys. A \textbf{45} (2012), 473001
doi:10.1088/1751-8113/45/47/473001
[arXiv:1205.5040 [hep-th]].


%\cite{Brigante:2008gz}
\bibitem{Brigante:2008gz}
M.~Brigante, H.~Liu, R.~C.~Myers, S.~Shenker and S.~Yaida,
%``The Viscosity Bound and Causality Violation,''
Phys. Rev. Lett. \textbf{100} (2008), 191601
doi:10.1103/PhysRevLett.100.191601
[arXiv:0802.3318 [hep-th]].

%\cite{Ge:2008ni}
\bibitem{Ge:2008ni}
X.~H.~Ge, Y.~Matsuo, F.~W.~Shu, S.~J.~Sin and T.~Tsukioka,
%``Viscosity Bound, Causality Violation and Instability with Stringy Correction and Charge,''
JHEP \textbf{10}, 009 (2008)
doi:10.1088/1126-6708/2008/10/009
[arXiv:0808.2354 [hep-th]].


%\cite{Ge:2009eh}
\bibitem{Ge:2009eh}
X.~H.~Ge and S.~J.~Sin,
%``Shear viscosity, instability and the upper bound of the Gauss-Bonnet coupling constant,''
JHEP \textbf{05}, 051 (2009)
doi:10.1088/1126-6708/2009/05/051
[arXiv:0903.2527 [hep-th]].






%\cite{Feng:2015oea}
\bibitem{Feng:2015oea}
X.~H.~Feng, H.~S.~Liu, H.~L\"u and C.~N.~Pope,
%``Black Hole Entropy and Viscosity Bound in Horndeski Gravity,''
JHEP \textbf{11} (2015), 176
doi:10.1007/JHEP11(2015)176
[arXiv:1509.07142 [hep-th]].



%\cite{Sadeghi:2020lfe}
\bibitem{Sadeghi:2020lfe}
M.~Sadeghi,
%``AdS black brane solution surrounded by quintessence in massive gravity and KSS bound,''
Mod. Phys. Lett. A \textbf{36} (2021) no.28, 2150202
doi:10.1142/S0217732321502023
[arXiv:2007.09688 [hep-th]].


%\cite{Bravo-Gaete:2020lzs}
\bibitem{Bravo-Gaete:2020lzs}
M.~Bravo-Gaete and F.~F.~Santos,
%``Complexity of four-dimensional hairy anti-de-Sitter black holes with a rotating string and shear viscosity in generalized scalar\textendash{}tensor theories,''
Eur. Phys. J. C \textbf{82}, no.2, 101 (2022)
doi:10.1140/epjc/s10052-022-10064-y
[arXiv:2010.10942 [hep-th]].


%\cite{Bravo-Gaete:2021hlc}
\bibitem{Bravo-Gaete:2021hlc}
M.~Bravo-Gaete and M.~M.~Stetsko,
%``Planar black holes configurations and shear viscosity in arbitrary dimensions with shift and reflection symmetric scalar-tensor theories,''
Phys. Rev. D \textbf{105}, no.2, 024038 (2022)
doi:10.1103/PhysRevD.105.024038
[arXiv:2111.10925 [hep-th]].


%\cite{Bravo-Gaete:2022lno}
\bibitem{Bravo-Gaete:2022lno}
M.~Bravo-Gaete, F.~F.~Santos and H.~Boschi-Filho,
%``Shear viscosity from black holes in generalized scalar-tensor theories in arbitrary dimensions,''
[arXiv:2201.07961 [hep-th]].

%\cite{Mamo:2012sy}
\bibitem{Mamo:2012sy}
K.~A.~Mamo,
%``Holographic RG flow of the shear viscosity to entropy density ratio in strongly coupled anisotropic plasma,''
JHEP \textbf{10}, 070 (2012)
doi:10.1007/JHEP10(2012)070
[arXiv:1205.1797 [hep-th]].


%\cite{SupernovaSearchTeam:1998fmf}
\bibitem{SupernovaSearchTeam:1998fmf}
A.~G.~Riess \textit{et al.} [Supernova Search Team],
%``Observational evidence from supernovae for an accelerating universe and a cosmological constant,''
Astron. J. \textbf{116} (1998), 1009-1038
doi:10.1086/300499
[arXiv:astro-ph/9805201 [astro-ph]].


%\cite{SupernovaCosmologyProject:1998vns}
\bibitem{SupernovaCosmologyProject:1998vns}
S.~Perlmutter \textit{et al.} [Supernova Cosmology Project],
%``Measurements of $\Omega$ and $\Lambda$ from 42 high redshift supernovae,''
Astrophys. J. \textbf{517}, 565-586 (1999)
doi:10.1086/307221
[arXiv:astro-ph/9812133 [astro-ph]].




%\cite{Planck:2013pxb}
\bibitem{Planck:2013pxb}
P.~A.~R.~Ade \textit{et al.} [Planck],
%``Planck 2013 results. XVI. Cosmological parameters,''
Astron. Astrophys. \textbf{571} (2014), A16
doi:10.1051/0004-6361/201321591
[arXiv:1303.5076 [astro-ph.CO]].


%\cite{Wang:1999fa}
\bibitem{Wang:1999fa}
L.~M.~Wang, R.~R.~Caldwell, J.~P.~Ostriker and P.~J.~Steinhardt,
%``Cosmic concordance and quintessence,''
Astrophys. J. \textbf{530} (2000), 17-35
doi:10.1086/308331
[arXiv:astro-ph/9901388 [astro-ph]].


%\cite{Tsujikawa:2013fta}
\bibitem{Tsujikawa:2013fta}
S.~Tsujikawa,
%``Quintessence: A Review,''
Class. Quant. Grav. \textbf{30} (2013), 214003
doi:10.1088/0264-9381/30/21/214003
[arXiv:1304.1961 [gr-qc]].


%\cite{Kiselev:2002dx}
\bibitem{Kiselev:2002dx}
V.~V.~Kiselev,
%``Quintessence and black holes,''
Class. Quant. Grav. \textbf{20} (2003), 1187-1198
doi:10.1088/0264-9381/20/6/310.


 %\cite{Letelier:1979ej}
\bibitem{Letelier:1979ej}
P.~S.~Letelier,
%``CLOUDS OF STRINGS IN GENERAL RELATIVITY,''
Phys. Rev. D \textbf{20} (1979), 1294-1302
doi:10.1103/PhysRevD.20.1294




%\cite{Ghosh:2014pga}
\bibitem{Ghosh:2014pga} 
S.~G.~Ghosh, U.~Papnoi and S.~D.~Maharaj,
%``Cloud of strings in third order Lovelock gravity,''
Phys.\ Rev.\ D {\bf 90}, no. 4, 044068 (2014)
doi:10.1103/PhysRevD.90.044068
[arXiv:1408.4611 [gr-qc]].


%\cite{Herscovich:2010vr}
\bibitem{Herscovich:2010vr}
E.~Herscovich and M.~G.~Richarte,
%``Black holes in Einstein-Gauss-Bonnet gravity with a string cloud background,''
Phys. Lett. B \textbf{689} (2010), 192-200
doi:10.1016/j.physletb.2010.04.065
[arXiv:1004.3754 [hep-th]].


%\cite{Lee:2014dha}
\bibitem{Lee:2014dha} 
T.~H.~Lee, D.~Baboolal and S.~G.~Ghosh,
%``Lovelock black holes in a string cloud background,''
Eur.\ Phys.\ J.\ C {\bf 75}, no. 7, 297 (2015)
doi:10.1140/epjc/s10052-015-3515-5
[arXiv:1409.2615 [gr-qc]].

%\cite{Mazharimousavi:2015sfo}
\bibitem{Mazharimousavi:2015sfo} 
S.~H.~Mazharimousavi and M.~Halilsoy,
%``Cloud of strings as source in $2+1$ -dimensional $f\left( R\right) =R^{n}$ gravity,''
Eur.\ Phys.\ J.\ C {\bf 76}, no. 2, 95 (2016)
doi:10.1140/epjc/s10052-016-3954-7
[arXiv:1511.00603 [gr-qc]].

%\cite{Graca:2016cbd}
\bibitem{Graca:2016cbd} 
J.~P.~Morais Graça, G.~I.~Salako and V.~B.~Bezerra,
%``Quasinormal modes of a black hole with a cloud of strings in Einstein-Gauss-Bonnet gravity,''
Int.\ J.\ Mod.\ Phys.\ D {\bf 26}, no. 10, 1750113 (2017)
doi:10.1142/S0218271817501139
[arXiv:1604.04734 [gr-qc]].

%\cite{Ranjbari:2019ktp}
\bibitem{Ranjbari:2019ktp}
H.~Ranjbari, M.~Sadeghi, M.~Ghanaatian and G.~Forozani,
%``Critical behavior of AdS Gauss\textendash{}Bonnet massive black holes in the presence of external string cloud,''
Eur. Phys. J. C \textbf{80}, no.1, 17 (2020)
doi:10.1140/epjc/s10052-019-7592-8
[arXiv:1911.10803 [hep-th]].

%\cite{Glavan:2019inb}
\bibitem{Glavan:2019inb}
D.~Glavan and C.~Lin,
%``Einstein-Gauss-Bonnet Gravity in Four-Dimensional Spacetime,''
Phys. Rev. Lett. \textbf{124}, no.8, 081301 (2020)
doi:10.1103/PhysRevLett.124.081301
[arXiv:1905.03601 [gr-qc]]. 

%\cite{Kiselev:2002dx}
\bibitem{Kiselev:2002dx}
V.~V.~Kiselev,
%``Quintessence and black holes,''
Class. Quant. Grav. \textbf{20}, 1187-1198 (2003)
doi:10.1088/0264-9381/20/6/310
[arXiv:gr-qc/0210040 [gr-qc]].



%\cite{Chen:2008ra}
\bibitem{Chen:2008ra}
S.~Chen, B.~Wang and R.~Su,
%``Hawking radiation in a $d$-dimensional static spherically-symmetric black Hole surrounded by quintessence,''
Phys. Rev. D \textbf{77} (2008), 124011
doi:10.1103/PhysRevD.77.124011
[arXiv:0801.2053 [gr-qc]].


   %\cite{Shepherd:2015dse}
\bibitem{Shepherd:2015dse} 
B.~L.~Shepherd and E.~Winstanley,
%``Dyons and dyonic black holes in ${\mathfrak {su}}(N)$ Einstein-Yang-Mills theory in anti-de Sitter spacetime,''
Phys.\ Rev.\ D {\bf 93}, no. 6, 064064 (2016)
doi:10.1103/PhysRevD.93.064064
[arXiv:1512.03010 [gr-qc]].
%%CITATION = doi:10.1103/PhysRevD.93.064064;%%

%\cite{Myers:2010ru}
\bibitem{Myers:2010ru}
R.~C.~Myers and B.~Robinson,
%``Black Holes in Quasi-topological Gravity,''
JHEP \textbf{08}, 067 (2010)
doi:10.1007/JHEP08(2010)067
[arXiv:1003.5357 [gr-qc]].


\bibitem{Ref26}
G.~Policastro, D.~T.~Son and A.~O.~Starinets,%``From AdS/CFT correspondence to hydrodynamics,''
JHEP {\bf 0209}, 043 (2002)
[hep-th/0205052].

\bibitem{Ref27}
P.~Kovtun, D.~T.~Son and A.~O.~Starinets,
%``Holography and hydrodynamics: Diffusion on stretched horizons,''
JHEP {\bf 0310}, 064 (2003)
[hep-th/0309213].


%\cite{Brigante:2007nu}
\bibitem{Brigante:2007nu}
M.~Brigante, H.~Liu, R.~C.~Myers, S.~Shenker and S.~Yaida,
%``Viscosity Bound Violation in Higher Derivative Gravity,''
Phys. Rev. D \textbf{77}, 126006 (2008)
doi:10.1103/PhysRevD.77.126006
[arXiv:0712.0805 [hep-th]].



%\cite{Ge:2020tid}
\bibitem{Ge:2020tid}
X.~H.~Ge and S.~J.~Sin,
%``Causality of black holes in 4-dimensional Einstein\textendash{}Gauss\textendash{}Bonnet\textendash{}Maxwell theory,''
Eur. Phys. J. C \textbf{80}, no.8, 695 (2020)
doi:10.1140/epjc/s10052-020-8288-9
[arXiv:2004.12191 [hep-th]].



%\cite{Sadeghi:2019muh}
\bibitem{Sadeghi:2019muh}
M.~Sadeghi and H.~Ranjbari,
%``Does a Cloud of Strings Affect Shear Viscosity Bound?,''
Class. Quant. Grav. \textbf{36}, no.20, 205012 (2019)
doi:10.1088/1361-6382/ab436e
[arXiv:1905.12856 [hep-th]].




 
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{hyperref}
\begin{document}
\title{\bf The effect of three matters on KSS bound }
\author{ %Shahrokh Parvizi$^1$\thanks{Corresponding author: Email:parvizi@modares.ac.ir}  \hspace{2mm} and
	Mehdi Sadeghi\thanks{Email:  mehdi.sadeghi@abru.ac.ir}\hspace{2mm}\\
	%	{\small {\em $^1$Department of Physics, School of Sciences,}}\\
	%   {\small {\em Tarbiat Modares University, P.O.Box 14155-4838, Tehran, Iran }}\\
	{\small {\em Department of Physics, School of Sciences,}}\\
	{\small {\em Ayatollah Boroujerdi University, Boroujerd, Iran}}
}
\date{\today}
\maketitle

\abstract{In this paper we introduce the black brane solutions in AdS space in 4-dimensional (4D) Einstein-Gauss-Bonnet-Yang-Mills theory in the presence of string cloud and quintessence. Shear viscosity to entropy density ratio is computed via fluid-gravity duality, as a transport coefficient for this model.}\\

\noindent PACS numbers: 11.10.Jj, 11.10.Wx, 11.15.Pg, 11.25.Tq\\
%\pacs{11.10.Jj, 11.10.Wx, 11.15.Pg, 11.25.Tq}

\noindent \textbf{Keywords:} AdS/CFT duality, Fluid-Gravity duality, Shear viscosity, Green-Kubo formula

%--------------------------------------------------------------------------
\section{Introduction} \label{intro}

\indent The AdS/CFT duality \cite{Maldacena}-\cite{Witten:1998qj} originates from string theory and provides a new perspective on quantum gravity. In its most general form, it is known as a gauge/gravity duality. This duality relates weakly coupled gravitational theories in $(D + 1)$-dimensional AdS space-time with strongly coupled conformal field theories (CFTs) defined on the $D$-dimensional boundary of AdS. In particular, it gives support to study the transport coefficients from hydrodynamics to the quark-gluon plasma formed at relativistic heavy-ion collisions \cite{Jiang:2017imk}-\cite{Hartnoll:2009sz}.  Shear viscosity $\eta$ \cite{D.T. Son}-\cite{Kovtun:2004de} as one of these coefficients is the most well-known coefficient calculated by this duality, especially fluid-gravity duality\cite{Bhattacharyya:2007vjd}-\cite{Bhattacharya:2011tra}.\\
The Kovtun-Son-Starinets (KSS)  bound states that the ratio $\eta/s$ has a lower bound,
$\frac{\eta }{s} \ge \frac{1 }{4\, \pi }$, for all relativistic quantum field theories \cite{Son:2007vk},\cite{Kovtun:2004de} and can be interpreted as the Heisenberg uncertainty principle \cite{Policastro:2001yc},\cite{Kovtun:2012rj}  where $\eta$ and s are the shear viscosity and entropy density, respectively. However, this conjecture violates for some theories like the Gauss-Bonnet  gravity \cite{Brigante:2008gz}, the Gauss-Bonnet gravity in presence of $U(1)$ gauge field \cite{Ge:2008ni},\cite{Ge:2009eh}, Horndeski theory \cite{Feng:2015oea}, massive gravity for $c_i<0$ \cite{Sadeghi:2020lfe}, scalar-tensor gravity\cite{Bravo-Gaete:2020lzs}-\cite{Bravo-Gaete:2022lno} and anisotropic black brane \cite{Mamo:2012sy}.\\
Observational cosmology shows the accelerated expansion of the Universe \cite{SupernovaSearchTeam:1998fmf},\cite{SupernovaCosmologyProject:1998vns}. Therefore, the Universe at present is becoming filled with a strange substance. This outcome is confirmed by the measurement of the Cosmic Microwave Background (CMB) using PLANCK space Satellite \cite{Planck:2013pxb}. It is believed that this expansion is due to a substantial negative pressure arising from the dark energy that is up to 70$\%$ of the total energy of the Universe. There are two sources of this negative pressure: one is cosmological constant and the other is the so called quintessence \cite{Wang:1999fa},\cite{Tsujikawa:2013fta} which acts as a repulsive force against the gravity. Quintessence is described by a scalar field and a parameter $\omega$, defined as the ratio of the pressure to the energy density of the dark energy. Black hole solution surrounded by the quintessence introduced in \cite{Kiselev:2002dx}.\\
String theory in which the fundamental ingredient of our universe is assumed to be one-dimensional strings, instead of particles, tries to merge the gravitational theory and quantum mechanics. An extension of this idea is to consider a cloud of strings \cite{Letelier:1979ej}-\cite{Ranjbari:2019ktp} and study its possible
measurable effects on long range gravitational fields of some
sources, shuch as black holes. Letelier \cite{Letelier:1979ej} has tried to extend the idea of a string to a black hole, possibly to reveal some properties of the blac khole. Ghosh and et al. obtained a generalization for third-order Lovelock gravity \cite{Ghosh:2014pga}, and Herscovich and et al. \cite{Herscovich:2010vr} obtained the solution for Einstein-Gauss-Bonnet theory in the Letelier spacetime.\\ 
General relativity (GR) provides the standard description of gravity and Gauss-Bonnet gravity is an extension gravity theory. Recently a new proposal has been offered wherein GB term is made  in $4D$ dimentions\cite{Glavan:2019inb}. In that ,the GB coupling is scaled as $\alpha \to \frac{\alpha}{D-4}$ and thereby canceling out $(D-4)$ factor in the equation, and then taking the limit $(D\to4)$. This results into an effective equation in $4D$ which is in fact the Einstein-Gauss-Bonnet (EGB)  equation written in $D=4$. Then it could be solved in spacetime with some specific symmetries for different situations, black holes and cosmology.\\
In this paper, we consider $4D$ Einstein-Gauss-Bonnet-Yang-Mills theory (EGBYM) in the presence of string cloud and quintessence. Then, by calculation of $\frac{\eta}{s}$ via fluid-gravity duality we study the effects of quintessence, string cloud and Yang-Mills charge on the field theory dual side.\\
%--------------------------------------------------------------------------
\section{$4D$ AdS Einstein-Gauss-Bonnet-Yang-Mills in presence of string cloud and quintessence black brane}
\label{sec2}

\indent The action of AdS Einstein-Gauss-Bonnet-Yang-Mills in the presence of string cloud and quintessence is:
\begin{equation}\label{Action}
I =\frac{1}{16\pi }\int{d^Dx\sqrt{-g}\Bigg[R-2 \Lambda+\frac{\alpha'}{D-4}\mathcal{G}+F^{(a)}_{\mu \alpha }F^{(a)\mu \alpha}\Bigg]}+S_{\text{cs}}+S_{\text{quint}},
\end{equation}
 where $ R $ is the scalar curvature, $\Lambda=\frac{-(D-1)(D-2)}{2 l^2}$ , $\alpha'$ is a (positive) Gauss-Bonnet coupling constant with dimension $(\text{length})^2$ , $  \mathcal{G}=R^2-4R_{\mu \nu}R^{\mu \nu}+R_{\mu \nu \rho \sigma }R^{\mu \nu \rho \sigma} $ , $F^{(a)}_{\mu \nu } =\partial _{\mu } A^{(a)}_{\nu } -\partial _{\nu } A^{(a)}_{\mu } -i[A^{(a)}_{\mu }, A^{(a)}_{\nu }]$ is the Cartan subalgebra of $SU(2) $ Yang-Mills  field strength tensor in which the gauge coupling constant is 1, $A_{\nu }$'s are the Cartan subalgebra of the $SU(2)$ gauge group Yang-Mills potentials.\\
%$S_{\text{quint}}$ ,the last term in (\ref{Action}), is the action for quintessence matter. A general action for quintessence in D-dimensional space-time is,
%\begin{equation}\label{Quin}
%S_{\text{quint}}= \frac{1}{2\kappa}\int{d^Dx\sqrt{-g}\Bigg[-\frac{1}{2}(\vec{\nabla} \phi)^2-V(\phi)\Bigg]},
%\end{equation}
%Because of simplicity and preserving space-time symmetry, quintessence is described by an interacting scalar field. 
By considering the following ansatz for the metric, 
\begin{equation}\label{metric1}
ds^{2} =-f(r)N^2(r)dt^{2} +\frac{dr^{2}}{f(r)} +r^2d\Omega^2_{2,\kappa},
\end{equation}where $\kappa = -1,0,1 $ and where  
$\Omega_{D-2}$ is the volume of unit $(D-2)$-sphere.\\
$f(r)$ is defined by a new variable $\phi(r)$,
\begin{equation}\label{f}
f(r)=\kappa-r^2\phi(r).
\end{equation}
The action of cloud of string is as follows,\\ 
\begin{equation}\label{CS}
S_{\text{cs}}=\int_{\Sigma} P \sqrt{-\gamma} d\tau d\sigma,
\end{equation}
where $P$ is non-negative constant related to string tension, $(\tau, \sigma)=(\lambda^0,\lambda^1)$ is a coordiantes of worldsheet of string with $\lambda^0$ and $\lambda^1$ are timelike and spacelike parameters
respectively, $\gamma_{ab}$ is the induced metric on the world sheet is as follows,
 \begin{equation}
\gamma_{ab}=g_{\mu \nu}(x) \frac{\partial x^{\mu}}{\partial \lambda^a}\frac{\partial x^{\nu}}{\partial \lambda^b},
 \end{equation}
and $\gamma=\text{det} \gamma_{ab}\nonumber$ is the determinant of the induced metric.\\ 
The bivector of string worldsheet $\Sigma$ is given by,\\
\begin{equation}
{\Sigma}^{\mu \nu}={\epsilon}^{a b} \frac{\partial x^{\mu}}{\partial \lambda^a}\frac{\partial x^{\nu}}{\partial \lambda^b},
\end{equation}
where ${\epsilon}^{a b}$ is the Levi-Civita tensor in two dimensions, $\epsilon^{01} = -\epsilon^{10} = 1$.\\
Therefore, the action of cloud of string (\ref{CS}) can be written as,
\begin{equation}
S_{\text{cs}}=P \int_{\Sigma}\sqrt{-\frac{1}{2}{\Sigma}^{\mu \nu} {\Sigma}_{\mu \nu}} d\tau d\sigma,
\end{equation}
The non-zero components of the bivector $\Sigma$ is $\Sigma^{tr}=-\Sigma^{rt}=-\frac{a}{P r^{D-2}}$. Where $a$ is an integration constant which is related to the cloud of string \cite{Letelier:1979ej}-\cite{Ranjbari:2019ktp}.\\
The action of quintessence is introduce by Kiselev \cite{Kiselev:2002dx} and written in terms of scalar field as follows,
\begin{equation}
S_{\text{quint}} =\frac{1}{16\pi G }\int{d^Dx\sqrt{-g}\Big[-\frac{1}{2}(\vec{\nabla} \phi)^2-V(\phi)\Big]}.
\end{equation}
The energy-momentum tensor of the quintessence dark energy in $D$ dimensions can be described by
\cite{Chen:2008ra} ,
\begin{align}
T_{t}\,^{t}&=T_{r}\,^{r}=\rho=\frac{\omega \alpha (D-1) (D-2) }{4 r^{(D-1)  (\omega +1)}},\\
T_{x_i}\,^{x_i}&=\frac{\rho}{D-2}((D-1) \omega+1),\,\, i=(1,...,D-2),
\end{align} 
where $\omega$ ($-1<\omega<-\frac{D-3}{D-1}$) and $\alpha$ are the quintessential state parameter and the positive normalization constant related to the density of quintessence, respectively.\\
By variation of the action with respect to $A^{(a)}_{\mu }$ we have, \\
 \begin{eqnarray}\label{EOM-YM}
\nabla_{\mu }F^{(a)\mu \nu } =0,
\end{eqnarray}
For solving Eq.(\ref{EOM-YM}), we consider the ansatz for the gauge field as follows\cite{Shepherd:2015dse},
\begin{equation}\label{background}
{\bf{A}}^{(a)} =\frac{i}{2}h(r)dt\begin{pmatrix}1 & 0 \\ 0 & -1\end{pmatrix},
\end{equation}
so,
\begin{equation}
h(r)=  C_2+Q\int^{r}\frac{ 1}{u^{D-2}}du,
\end{equation}
\begin{equation}
h'(r)= \frac{ Q}{r^{D-2}},
\end{equation}
By plugging the  ansatz \ref{metric1} into the action  Eq.(\ref{Action}) yields,
\begin{align}\label{EOMI}
 &I =\frac{\Omega_{D-2}(D-2)}{16\pi }\nonumber\\&\int{dt dr N(r)\Bigg[r^{D-1}\phi\bigg(1+\alpha'(D-3)\phi\bigg)+\frac{r^{D-1}}{l^2}+\frac{2Q^2 r^{3-D}}{(D-3)(D-2)}-\frac{2 a r}{(D-2)}-\frac{\alpha}{r^{\omega (D-1)}} \Bigg]'},
\end{align}
in which $'$ denotes derivative with respect to $r$. Equation of motion is given by variation
of $\delta N(r)$ \cite{Myers:2010ru}. The function $ \phi $ is given by solving for the roots of a quadratic polynomial,
\begin{equation}\label{phi}
\phi+\alpha'(D-3)\phi^2=\frac{16 \pi M}{(D-2)r^{D-1}\Omega_{D-2}}-\frac{1}{l^2}-\frac{2Q^2 r^{4-2D}}{(D-3)(D-2)}-\frac{2a}{(D-2)r^{D-2}}+\frac{\alpha}{r^{(\omega+1) (D-1)}},
\end{equation}
where $M$ is a constant.\\
By solving Eq.(\ref{phi}) and plugging $\phi$ into Eq.(\ref{f}) we will have,
\begin{equation}
\begin{split}
& f(r)=\kappa-\frac{r^2}{2\alpha'(D-3)}\Bigg[-1\pm \\ &\sqrt{1-4(D-3)\alpha'\bigg(\frac{1}{l^2}+\frac{2Q^2 r^{4-2D}}{(D-3)(D-2)}-\frac{16 \pi M r^{1-D}}{(D-2)\Omega_{D-2}}+\frac{2 a r^{2-D} }{D-2}+\frac{\alpha}{r^{(D-1)(\omega+1)}}\bigg)}\Bigg].
\end{split}
\end{equation}
 The line elements of planar black branes can be written as,\\
\begin{equation}\label{metric}
ds^{2} =-H(r)N(r)^2dt^{2} +\frac{ dr^2 }{H(r)} +\frac{r^2}{l^2} \sum_{i=1}^{D-2}dx_i^2.
\end{equation}
That's easy to show $ N(r) $ is constant by variation of $ \phi(r) $ from Eq.(\ref{EOMI}).
\begin{equation}\label{H}
\begin{split}
	& H(r)=\frac{r^2}{2\alpha'(D-3)}\Bigg[1- \\ &\sqrt{1-4(D-3)\alpha'\bigg(\frac{1}{l^2}+\frac{2Q^2 r^{4-2D}}{(D-3)(D-2)}-\frac{16 \pi M r^{1-D}}{(D-2)\Omega_{D-2}}+\frac{2 a r^{2-D} }{D-2}+\frac{\alpha}{r^{(D-1)(\omega+1)}}\bigg)}\Bigg],
\end{split}
\end{equation}
where $M$ and $Q$ are integration constants proportional to the mass and charge of the black hole
respectively given by the following formulas:\\
 \begin{equation}
 \begin{split}
&M=\frac{(D-2)\Omega_{D-2}}{16 \pi}m,\\
&Q^2=\frac{(D-2)(D-3)}{2}q^2.\\
\end{split}
 \end{equation}
By defining new parameters as below,\\
 \begin{equation}
\begin{split}
&\frac{(D-3)\alpha'}{l^2}=\lambda_{gb},\\
&\frac{2 a  }{D-2}=A,
\end{split}
\end{equation} 
 
and substituting them into the Eq.(\ref{H}) we will have,
\begin{equation}\label{H1}
H(r)=\frac{r^2}{2\lambda_{gb} l^2}\Bigg[1- \sqrt{1-4\lambda_{gb} \bigg(1+\frac{q^2l^2}{r^{2D-4}}-\frac{ m l^2}{r^{D-1}}+\frac{A l^2}{r^{D-2}}+\frac{\alpha l^2}{r^{(D-1)(\omega+1)}}\bigg)}\Bigg].
\end{equation}
Black hole has an event horizon, so by applying the condition $H(r_+)=0$ we can find $m$ as follows,
\begin{equation}
m=\frac{r_+^{D-1}}{l^{2}}+\frac{q^2}{r_+^{D-3}}+A r_++\frac{\alpha l^2}{r_+^{(D-1)\omega}}.
\end{equation}
By plugging $m$ in Eq.(\ref{H1})
\begin{equation}
H(r)=\frac{r^2}{2\lambda_{gb} l^2}\Bigg[1-\sqrt{1-4\lambda_{gb}\Bigg(1-(\frac{r_+}{r})^{D-1}+l^2 q^2 \bigg(\frac{1}{r^{2D-4}}-(\frac{r_+}{r})^{D-1}\frac{1}{r_+^{2D-4}}\bigg)+\mathcal{P}\bigg)\Bigg)}\Bigg],
\end{equation}
where,
\begin{equation}
\mathcal{P}=A l^2\bigg(\frac{1 }{r^{D-2}}-(\frac{r_+}{r})^{D-1}\frac{1}{r_+^{D-2}}\bigg)+\frac{\alpha l^2}{r^{(D-1)}}\bigg(\frac{1}{r^{(D-1)\omega}}-\frac{1}{r_+^{(D-1)\omega}}\bigg).
\end{equation}
In AdS/CFT correspondence, the speed of light in the boundary CFT is simply $c=1$. So that, for the black brane solution in the asymptotic region we have $\lim_{r \to 0}{N(r)^2 H(r)}=1$ to recover a causal boundary. By applying this criterion we will have,\\
\begin{equation}
N^2=\frac{1+\sqrt{1-4\lambda_{gb}}}{2}.
\end{equation}
The temperature and the Hawking-Bekenstein entropy density follow as,
\begin{eqnarray}
&T=\frac{1}{4\pi \sqrt{g_{rr} g_{tt} } } \partial_r g_{tt}|_{r=r_+ } =\frac{Nr_{+}}{4\pi l^2}\Big[(D-1)+\frac{Al^2}{r_+^{D-2}}+\frac{(D-3)l^2q^2}{r_+^{2D-4}}-\frac{(D-1)\alpha \omega}{r_+^{(D-1)(\omega+1)}}\Big],\\
&s=\frac{4\pi}{V}\int d^{3}x \sqrt{-g}=4\pi\left(\frac{r_+}{l}\right)^{3}.
\end{eqnarray}

%--------------------------------------------------------------------------

\section{$\frac{\eta}{s}$ of this solution}
 \label{sec4}
By introducing new variables $ z=\frac{r}{r_+} $ , $ \omega=\frac{l^2}{r_+}\tilde{\omega} $ , $ k_3=\frac{l^2}{r_+ ^2}\tilde{k}_3 $, , $ \tilde{H}(z)=\frac{l^2}{r_+ ^2}H $, The line element in terms of new variables is as follows,
\begin{equation}\label{metricGB3}
ds^{2} =-N^2\tilde{H} dt^{2} +\frac{ dz^2 }{r_+^2\tilde{H}} +\frac{r_+^2 z^2}{l^2} \sum_{i=1}^{D-2}dx_i^2,
\end{equation}
where
\begin{equation}
\tilde{H}(z)=\frac{z^2}{2\lambda_{gb} }\Bigg[1-\sqrt{1-4\lambda_{gb}\Bigg(1-\frac{1}{z^{D-1}}+\frac{l^2 q^2}{r_+^{2D-4}} \bigg(\frac{1}{z^{2D-4}}-\frac{1}{z^{D-1}}\bigg)+\mathcal{P}\Bigg)}\Bigg],
\end{equation}
\begin{equation}
\mathcal{P}=\frac{A l^2}{r_+^{D-2}}\bigg(\frac{1 }{z^{D-2}}-\frac{1 }{z^{D-1}}\bigg)+\frac{\alpha l^2}{r_+^{(D-1)(\omega+1)}}\bigg(\frac{1}{z^{(D-1)\omega}}-\frac{1}{z^{(D-1)(\omega+1)}}\bigg).
\end{equation}
To calculate the shear viscosity, we perturb the background metric by $ h^{y}_{x} $. We call the perturbed part of metric  $\psi=h^{y}_{x}$ \cite{Ref26,Ref27}. 
\begin{equation}\label{metricGBYMCS}
ds^{2} =-N^2\tilde{H} dt^{2} +\frac{ dz^2 }{u^2\tilde{H}} +\frac{ z^2}{u^2 l^2}\Bigg( \sum_{i=1}^{D-2}dx_i^2+2\psi(t,\vec{x},z)dx_1dx_2\Bigg),
\end{equation}
where $u=\frac{1}{r_+^2}$.\\
Using Fourier decomposition, 
\begin{equation}
\psi(t,x_i,z)=\int{\frac{d^{D-1}k}{(2\pi)^{D-1}}e^{-i\tilde{\omega}t+\tilde{k_i} x_i}}\psi(k,z).
\end{equation}
We substitute the perturbed metric Eq.(\ref{metricGBYMCS}) in Eq.(\ref{Action}) and expand the action up to second order of $\psi$. Finally, the equation of motion for $\psi(z)$ can be obtained by variation of  the perturbed action with respect to $\psi$ as follows,\\
\begin{equation}\label{Mod}
(K \psi')'+\omega^2K_1\psi-k_i^2K_2\psi=0,
\end{equation}
where
 \begin{align}\label{54}
  &K=\frac{1}{16 \pi}\sqrt{-g}g^{zz}g^{x_i x_i}K_3(z),\nonumber\\
  &K_1=-\frac{1}{16 \pi}\sqrt{-g}g^{tt}K_3(z), \nonumber\\
  &K_2=\sqrt{-g}g^{x_i x_i}K_4(z), \nonumber\\
  &K_3(z)=1-\frac{2\lambda_{gb}}{D-3}[z^{-1} \tilde{H}'+z^{-2} (D-5) \tilde{H}], \nonumber\\
   &K_4(z)=1-\frac{2\lambda_{gb}}{(D-3)(D-4)}[\tilde{H}''+ (D-5)(D-6)z^{-2} \tilde{H}+2(D-5)z^{-1}\tilde{H}'], \nonumber\\
   \end{align} 
in which $'$ denotes derivative with respect to $z$. The factors $(D-5)$ and
$(D-6)$ in the expression of $K_4(z)$ is for $D>5$ of the Gauss-Bonnet theory.\\
The Green function is,
\begin{equation} 
G_{x_i x_j, x_i x_j}=\frac{K\psi'}{\psi}\, .
\end{equation}
The shear viscosity is calculated by Green-Kubo formula as the following,
\begin{equation}
\eta_{x_i x_j, x_i x_j}=\frac{-G_{x_i x_j, x_i x_j}}{i \omega},
\end{equation}
Now we write Eq.(\ref{Mod}) in terms of the shear viscosity, 
\begin{equation}
\partial_{z}\eta_{x_i x_j, x_i x_j}=(\frac{\eta^2_{x_i x_j, x_i x_j}}{K}-K_1)+\frac{i}{\omega}K_2k_i^2,
\end{equation}
and then we can compute the shear viscosity by requiring horizon regularity as follows,
\begin{equation}
\eta_{x_i x_j, x_i x_j}=\sqrt{K K_1}\Bigg|_{z=1}=\frac{1}{16\pi}\frac{r_+^{D-2}}{l^{D-2}}  \left(1-\frac{2 \lambda  }{D-3}\tilde{H}'(1)\right),
\end{equation}
The ratio of the shear viscosity to the entropy density for 4D charged black hole solutions
in Gauss-Bonnet gravity in the presence of string cloud and quintessence is then\\
\begin{equation}
\frac{\eta}{s}=\frac{1}{4\pi }[1-\frac{2 \lambda  }{D-3}\tilde{H}'(1)],
\end{equation}
\begin{equation}
\frac{\eta}{s}=\frac{1}{4\pi }\Bigg[1-\frac{2 \lambda  }{D-3}\Bigg((D-1)-(D-3) l^2 q^2 r_+^{4-2 D}+A l^2 r_+^{2-D}+\alpha l^2 (D-1) r_+^{(1-D)(1+\omega)}\Bigg)\Bigg].
\end{equation}
For $\lambda \to 0$ this value is $\frac{\eta}{s}=\frac{1}{4\pi }$. It means that the KSS bound is saturated for Einstein-Hilbert gravity.\\
For $q=A=\alpha=0$ we will have,
\begin{equation}
\frac{\eta}{s}=\frac{1}{4\pi }\big[1-2\frac{(D-1)}{(D-3)}\lambda \big],
\end{equation}
which is consistent with the literature\cite{Brigante:2007nu}.\\
In the limit $D \to 4$, we obtain
\begin{equation}
\frac{\eta}{s}=\frac{1}{4\pi }\Bigg[1-2 \lambda \Bigg(3- l^2 q^2 r_+^{-4}+A l^2 r_+^{-2}+3\alpha l^2  r_+^{-3(1+\omega)}\Bigg)\Bigg].
\end{equation}
For $A=\alpha=0$ the result is the same as \cite{Ge:2020tid},
\begin{equation}
\frac{\eta}{s}=\frac{1}{4\pi }\Bigg[1-2 \lambda \Big(3- l^2 q^2 r_+^{-4}\Big)\Bigg].
\end{equation}
%--------------------------------------------------------------------------
\section{Conclusion}

\noindent In summary, we have investigated the effect of string cloud , quintessence and Yang-Mills charge in $4D$ Gauss-Bonnet gravity on the field theory dual side by the fluid-gravity duality. Our result shows that theses three matters are affected in dual of Gauss-Bonnet gravity but they have no contribution in dual of Einstein-Hilbert gravity.   Therefore,  the KSS conjecture \cite{Kovtun:2004de} is saturated in the limit $\lambda \to 0$ . This conjecture tells us that the ratio $\eta/s$ has a lower bound, $\frac{\eta }{s} \ge \frac{\hbar }{4\, \pi \, k_{B} } $, for all relativistic quantum field theories at finite temperature without chemical potential \cite{Policastro:2001yc} and can be interpreted as the Heisenberg uncertainty principle \cite{Son:2007vk}. String cloud \cite{Sadeghi:2019muh} and quintessence \cite{Sadeghi:2020lfe} in Einstein massive gravity does not contribute to the $\frac{\eta}{s}$. However, this conjecture violates for higher derivative theories of gravity like the AdS Einstein-Gauss-Bonnet-Yang-Mills in the presence of string cloud and quintessence as we confirmed in section 3 . Our outcome shows that the coupling of this model is scaled as $\lambda \to \lambda \Big((D-1)-(D-3) l^2 q^2 r_+^{4-2 D}+A l^2 r_+^{2-D}+\alpha l^2 (D-1) r_+^{(1-D)(1+\omega)}\Big)$.
%--------------------------------------------------------------------------

\vspace{1cm}
\noindent {\large {\bf Acknowledgment} }  Author would like to thank Xian-Hui Ge , M.H. Abbasi ,and the referees of MPLA for useful comments and suggestions.

%--------------------------------------------------------------------------

\vspace{1cm}
\noindent {\large {\bf Data Availability } } This manuscript has no associated data.

%--------------------------------------------------------------------------

\end{document}